OpenStudy (anonymous):

Topic: Combination in a class of 10 boys and 15 girls ,in how many ways may a committee be formed out of 4 boys and 3 girls?

OpenStudy (raden):

10C4 multiply by 15C3

OpenStudy (anonymous):

thanks

OpenStudy (raden):

yw

OpenStudy (anonymous):

i will do it now w8

OpenStudy (anonymous):

95550?

OpenStudy (raden):

10C4 = 10!/(10-4)!4! = 10!/(6!4!) = 10.9.8.7.6!/(6!.4.3.2.1) = 210 15C3 = 15!/(15-3)!3! = 15!/(12!3!) = 15.14.13.12!/(12!.3.2.1) = 455 multiply of both = 210 * 455 = 95550 U R CORRECT

OpenStudy (anonymous):

thank u sir

OpenStudy (raden):

you're welcome